In a random sample of 64 trout, the average weight is 3 pounds, with standard deviation s = .5 pounds Find:

sample mean

standard error of the sample mean

margin of error 95%

95% confidence interval for the population mean, µ:
Answer #1


Solution :

Given that,

\bar x = 3

s = 0.5

n = 64

The standard error = (s /\sqrtn)

= (0.5 / \sqrt 64)

=0.06

The standard error =0.06

Degrees of freedom = df = n - 1 = 64 - 1 = 63

At 95% confidence level the t is ,

\alpha  = 1 - 95% = 1 - 0.95 = 0.05

\alpha / 2 = 0.05 / 2 = 0.025

t\alpha /2,df = t0.025,63 =1.998

Margin of error = E = t\alpha/2,df * (s /\sqrtn)

= 1.998 * (0.5 / \sqrt 64)

= 0.12

Margin of error =0.12

The 95% confidence interval estimate of the population mean is,

\bar x - E < \mu < \bar x + E

3 - 0.12 < \mu < 3 + 0.12

2.88 < \mu < 3.12

(2.88, 3.12 )